Math 550 (Section 1)  University of South Carolina
D. Meade  Fall 1996

Homework Problems for Chapter 4 Vector-Valued Functions

Section Page Problems
4.1  254 # 2, 3^{∗}, 10, 11, 13^{∗}, 17, 19(ad^{∗}), 20
4.2  262 # 1, 7^{∗}, 9, 10, 12, 13^{∗}
4.3  272 # 2, 3, 4^{∗}, 7^{∗}, 10, 15, 17(a^{∗})
4.4  286 # 3, 5, 7^{∗}, 14, 15^{∗}, 22, 25^{∗}, 27, 29^{∗}, 30
Review 289 # 1^{∗}, 9, 11^{∗}, 13, 23^{∗}, 32

\underline{Notes}

• Problems marked with an asterisk (^{∗}) will be collected and graded. In general, problems will be due the Monday following the lecture in which the section was discussed.

• These problems have been selected as the minimum amount of work that you should be doing for this course. Some problems will be straightforward computations; some will require the construction of a logical argument using the terminology and concepts discussed in class; some problems will require you to expand your knowledge (with given information) and apply this to a new setting.

• If you need more practice, there are many other problems in the text that are similar to the ones I have assigned.

• Be sure you know how to do all of these problems. Do not hesitate to ask questions – in class, by e-mail, or during office hours.